Energy required to move a body of mass m from an orbit of radius 2R to 3R is
(a) GMm/12R2
(b) GMm/3R2
(c) GMm/8R
(d) GMm/6R

Grade:6

3 Answers

Deepak Patra
9 years ago

(d) Energy required = (Potential energy of the Earth – mass system when mass is at distance 3R) – (Potential energy of the Earth-mass system when mass is at distance 2R)

= - GM/3R – (-GM/2R) = - GM/3R + GM/2R

= -2GMm+3GMm / 6R = GMm/6R
Mallikarjun
6 years ago
Potential energy U=-GMm/rTo get energy required to move a body of mass m from 2R to 3R we have to subtract the energy required to move the body from earth to 2R, from the energy required to move the body from earth to 3RHence U=-GMm/3R -(-GMm/2R) =GMm6R
Yash Chourasiya
3 years ago
Dear Student

Energyrequired = (Finalpotential – Initialpotential)m
Energyrequired = (−GMm/​3R)− (−GMm/2R)
​Energy Required = GMm/6R​

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya